OpenStudy (anonymous):

How do you expand (x-3)^5 using the binomial theorem?

OpenStudy (anonymous):

Binomial theorem: \[(a+b)^n=\sum_{k=0}^n\begin{pmatrix}n\\k\end{pmatrix}a^{n-k}b^k\]

OpenStudy (anonymous):

Where \(\begin{pmatrix}n\\k\end{pmatrix}=\dfrac{n!}{k!(n-k)!}\).

OpenStudy (anonymous):

i dont understand that equation

OpenStudy (anonymous):

So for this problem, you have \[\begin{align*}(x-3)^5&=\sum_{k=0}^5\begin{pmatrix}5\\k\end{pmatrix}x^{5-k}(-3)^k\\ &=\begin{pmatrix}5\\0\end{pmatrix}x^{5-0}(-3)^0+\begin{pmatrix}5\\1\end{pmatrix}x^{5-1}(-3)^1+\begin{pmatrix}5\\2\end{pmatrix}x^{5-2}(-3)^2\\ &~~~~+\begin{pmatrix}5\\3\end{pmatrix}x^{5-3}(-3)^3+\begin{pmatrix}5\\4\end{pmatrix}x^{5-4}(-3)^4+\begin{pmatrix}5\\5\end{pmatrix}x^{5-5}(-3)^5 \\&=\cdots\end{align*}\]

OpenStudy (anonymous):

so would my final answer be x^5+x^4+x^3+x^2+x-183?

OpenStudy (anonymous):

No. You're forgetting the coefficients. Are you familiar with the binomial coefficient?

OpenStudy (anonymous):

no I dont know what that is?

OpenStudy (anonymous):

is it 1, 5 10 10 5 1?

OpenStudy (anonymous):

Yes! Also, you have to keep track of the \((-3)^{\cdots}\). They also contribute to the coefficients of each power of x.

OpenStudy (anonymous):

For example, I'll work out the third term: \[\begin{pmatrix}5\\2\end{pmatrix}x^{5-2}(-3)^2=5x^3(-9)=-45x^3\]

OpenStudy (anonymous):

ok is it x^5+5x^4+10x^3+10x^2+5x+24267?

OpenStudy (anonymous):

No, you're still missing the powers of (-3) in each term.

OpenStudy (anonymous):

Would the first term is x^5 (-3^0) right? and the second would be x^4 (-3^1)*5? and so on

OpenStudy (anonymous):

So would the final answer be x^5-15x^4+90x^3-270x^2+405x-243?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

YESS Thankyou very much

OpenStudy (anonymous):

You're welcome